OpenStudy (anonymous):

The legs of a right triangle measure 8ft and 15ft, what is the length of the hypotenuse?

hartnn (hartnn):

do u know pythagoras theorem ?

OpenStudy (ryan123345):

\[A^2+B^2=C^2\]

OpenStudy (anonymous):

No

OpenStudy (ryan123345):

Plug in A and B for the legs and C for the hypotenuse.

OpenStudy (ryan123345):

\[8^2+15^2=C^2\]

hartnn (hartnn):

 \(a^2+b^2=c^2\) using pythagoras theorem.